# Calibrating Bayesian Learning via Regularization, Confidence Minimization, and Selective Inference

### v2

@@ V-A Setting and Metrics
visualize the bins that have number of samples no smaller than 100 to avoid undesired statistical noise [55, 56]); (ii) the ECE, which evaluates the average discrepancy between per-bin accuracy and per-bin confidence as defined in (6); (iii) ID accuracy, marked as “Acc”, which is the proportion of the ID data for which a correct prediction is made; (iv) OOD detection probability, which is the optimal probability of successfully detecting OOD data, as defined in (20); and (v) ID coverage rate, which is the fraction of selected data points on the ID test data set for selective calibration. We refer to Appendix -B for all the experimental details [^1] Code can be found at https://github.com/kclip/Calibrating-Bayesian-Learning..
@@ V-D Can Selective Calibration Compensate for the ID Performance Loss of OCM?
So far, we have seen that there is generally a trade-off between ID and OOD performance. In particular, from Fig. 9, it was concluded that, for a given fixed level of OOD detection probability, calibration-regularized learning improved ID calibration at the cost of the ID accuracy for BNN-OCM. In this subsection, we ask whether selective calibration can ensure a synergistic use of OCM and calibration-aware regularization, guaranteeing that CBNN-OCM achieves the best ID and OOD performance levels. To address this question, we vary the ID coverage rate, i.e., the fraction of accepted ID samples, by changing the hyperparameter $\tau$ in (IV-C). The resulting ID and OOD performance is shown in Fig. 10 for SBNN-OCM and SCBNN-OCM.
@@ VI Conclusion
In this paper, we have proposed SCBNN-OCM, a general framework that generalizes variational inference-based Bayesian learning to target both ID and OOD calibration. To improve ID calibration, we have introduced a regularizer based on the calibration error, while OOD calibration is enhanced by means of data augmentation based on confidence minimization. In order to facilitate the synergistic use of calibration-aware regularization and OCM regularization, we have finally introduced a selective calibration strategy that rejects examples that are likely to have poor ID calibration performance. Numerical results have illustrated the challenges in ensuring both ID and OOD performance, as schemes designed for ID calibration may end up hurting OOD calibration, and vice versa. That said, thanks to selective calibration, SCBNN-OCM was observed to achieve the best ID and OOD performance as compared to the other benchmarks as long as one allows for a sufficiently large number of rejected samples. For instance, as compared to standard FNN, SCBNN-OCM achieves a $25\%$ improvement in accuracy, a $20\%$ drop in ECE, and a $50\%$ improvement in OOD detection probability, at the cost of reduced ID coverage rate of around $50\%$. Interesting directions for future work include extending the proposed framework to likelihood-free, simulation-based inference [57], to continual learning [58], and to large language models [13].
-A Non-Parametric Outlier Score Vectors
In this subsection, we specify the details for non-parametric outlier score vector $s(x|\theta^{\text{FNN}})$ described in Sec. IV-B2, which can be implemented by following the sklearn package in Python. As mentioned in the text, the first entry is given by (35) with the Gaussian kernel $\kappa\left(||z-z^{\text{tr}}_{i}||\right)=\exp\left(-||z-z_{i}^{\text{tr}}||^{2}/h\right)$ given the bandwidth parameter $h>0$.
@@
$s_{3}(x|\theta^{\text{FNN}})=\ \sum_{i=1}^{|\mathcal{X}^{\text{tr}}|}\alpha_{i}\kappa\left(||z-z^{\text{tr}}_{i}||\right)-\rho,$ (43)
in which the coefficients $\alpha_{i}\geq 0$ for $i=1,...,|\mathcal{X}^{\text{tr}}|$, satisfying $\sum_{i=1}^{|\mathcal{X}^{\text{tr}}|}\alpha_{i}=1$, are optimized along with the offset parameter $\rho\in\mathbb{R}$ in order to separate the training data $\{z_{i}^{\text{tr}}\}_{i=1}^{|\mathcal{X}^{\text{tr}}|}$ from the origin in the feature space associated with the kernel $\kappa(\cdot)$ [60, Sec. 5].
@@
$s_{4}(x|\theta^{\text{FNN}})=||z-z_{i}^{\text{tr}}||_{(k)},$ (44)
where $||z-z_{i}^{\text{tr}}||_{(k)}$ is the $k$-th smallest value in the set of distances $\{||z-z_{i}^{\text{tr}}||\}_{i=1}^{|\mathcal{X}^{\text{tr}}|}$.
-B Experiment Details
@@
-B1 Architecture and Training Details for Calibration-Regularized Learning
@@
Hyperparameters: For fair comparison, we use the same training policy for both frequentist and Bayesian learning. Specifically, we use the SGD optimizer with momentum factor $0.9$ and train the model during $100$ epochs. We also decrease the learning rate by dividing its value by a factor of $5$ for every $30$ epochs with initial learning rate $0.1$ in a manner similar to [11]. The minibatch size used for single SGD update is set to $128$. For Bayesian learning, we set as prior $p(\theta)$ the Gaussian distribution that has zero-mean vector with diagonal covariance matrix having each element as $0.001$. The hyperparamter $\beta$ for the free energy (11) is set to $0.00035$. The ensemble size for Bayesian learning during training is set to $1$, while the ensemble size during testing is set to $20$. In a manner similar to [6], the calibration-based regularizer $\lambda$ in (9) and (17) is chosen as the value in set $\{0.2,0.4,0.6,0.8,1.0,2.0,3.0,...,10.0\}$ that achieves the lowest ECE, while preserving the accuracy drop no larger than $1.5\%$ as compared to a setting with $\lambda=0$. The corresponding accuracy and ECE are evaluated based on the validation data set $\mathcal{D}^{\text{val}}$, and the hyparparameter $\lambda$ for CFNN and for CBNN are chosen independently, which result in the values of $4$ and $0.8$ respectively in our experiments. With these choices, CFNN and CBNN can achieve lower ECE with acceptable accuracy drops. We use weighted MMCE regularizer for calibration-aware regularization term $\mathcal{E}(\theta|\mathcal{D}^{\text{tr}})$, which is defined as [4]
@@
where the kernel function is $\kappa(r_{i},r_{j})=\exp(-||r_{i}-r_{j}||/0.4)$, and $n_{c}$ is the number of correct examples, i.e., $n_{c}=\sum_{i=1}^{|\mathcal{D}^{\text{tr}}|}c_{i}.$
@@
-B2 Architecture and Training Details for OOD Detection
@@
Hyperparameters: Following the original OCM paper [11], OOD confidence minimization (21) and (III-C) is carried out by fine-tuning based on the corresponding pre-trained models, e.g., CBNN-OCM is obtained via fine-tuning with the OCM-regularized training loss given the pre-trained CBNN. Uncertainty data set $\mathcal{D}^{\text{u}}$ is constructed by randomly choosing $6,000$ input data from TinyImageNet. During fine-tuning, SGD optimizer with momentum factor of $0.9$ is adopted during $10$ epochs, each consisting 282 iterations. The initial learning rate is set to $0.001$, and we update the learning rate for each iteration by following [63]. Specifically, during SGD training, for each epoch, we first sample 9,000 examples from $\mathcal{D}^{\text{tr}}$ and evaluate the ID-related training measures for each SGD update via sampling 32 examples (minibatch size being 32) without replacement among the 9,000 examples; for the OOD-related training measures, we sample 64 examples (minibatch size being 64) among $\mathcal{D}^{\text{u}}$ with replacement. The hyperparameter of OCM regularizer $\gamma$ in (21) and (III-C) is chosen as in reference [11], i.e, $\gamma=0.5$, which increases OOD detection probability nearly to $1$ for all OCM schemes. Other hyperparameters are the same as calibration-regularized learning described in the previous subsection.
@@
-B3 Architecture and Training Details for Selective Calibration
@@
Hyperparameters: For selector training (IV-B1) and (IV-C), we use Adam optimizer, with learning rate $0.001$ and weight decay coefficient $10^{-5}$. We train the model for $5$ epochs, each epoch consisting $50,000$ iterations, and each iteration samples $32$ examples (minibatch size being $32$) among $\mathcal{D}^{\text{val}}$. Since the proposed SCBNN-OCM is more robust to the choives of $\eta$, the hyperparameter $\eta$ in (IV-B1) and (IV-C) is set to $0.01$ as reported in [12], and we set the kernel function in (LABEL:eq:sel-cal-loss-bayes) as $\kappa(r_{i},r_{j})=\exp(-||r_{i}-r_{j}||/0.2)$. We use the same ensemble size for training and testing as in calibration-regularized learning.

### v3

@@ V-A Setting and Metrics
visualize the bins that have number of samples no smaller than 100 to avoid undesired statistical noise [52]); (ii) the ECE, which evaluates the average discrepancy between per-bin accuracy and per-bin confidence as defined in (6); (iii) ID accuracy, marked as “Acc”, which is the proportion of the ID data for which a correct prediction is made; (iv) OOD detection probability, which is the optimal probability of successfully detecting OOD data, as defined in (20); and (v) ID coverage rate, which is the fraction of selected data points on the ID test data set for selective calibration. We refer to Appendix in supplementary document for all the experimental details [^1] Code can be found at https://github.com/kclip/Calibrating-Bayesian-Learning..
@@ V-D Can Selective Calibration Compensate for the ID Performance Loss of OCM?
So far, we have seen that there is generally a trade-off between ID and OOD performance. In particular, from Fig. 9, it was concluded that, for a given fixed level of OOD detection probability, calibration-regularized learning improved ID calibration at the cost of the ID accuracy for BNN-OCM. In this subsection, we ask whether selective calibration can ensure a synergistic use of OCM and calibration-aware regularization, guaranteeing that CBNN-OCM achieves the best ID and OOD performance levels. To address this question, we vary the ID coverage rate, i.e., the fraction of accepted ID samples, by changing the hyperparameter $\tau$ in (41). The resulting ID and OOD performance is shown in Fig. 16 for SBNN-OCM and SCBNN-OCM.
@@
V-E Memory Requirements and Computational Cost
Finally, we analyze the memory requirements and computational cost of the different schemes studied in this paper. To this end, we focus on the setting adopted for the experiments reported in Fig. 6, Fig. 8, and Fig. 16. We evaluate the number of model parameters and the number of floating-point operations (FLOPs) during inference. The number of FLOPs is reported normalized by the size of the ensemble, and is evaluated using a standard profiler [53]. The results, reported in Table II, can be used to determine the trade-off between the performance improvements observed in Fig. 6, Fig. 8, and Fig. 16 and the memory and computational costs. The overall number of FLOPs can be obtained from Table II by multiplying by the ensemble size, which is set to $20$ for inference and to $1$ for training. Furthermore, the number of training FLOPs per iteration must account for the cost of gradient evaluation, which is typically approximated as 22-55 times the inference cost [54, 55].
TABLE II: Number of parameters and number of inference FLOPs per ensemble size (FLOPs are calculated with image size $32\times 32$).
Model Parameters FLOPs/ensemble size
FNN 2.25M 0.3G
CFNN 2.25M 0.3G
BNN 4.5M 0.34G
CBNN 4.5M 0.34G
FNN-OCM 2.25M 0.3G
CFNN-OCM 2.25M 0.3G
BNN-OCM 4.5M 0.34G
CBNN-OCM 4.5M 0.34G
Selector 0.005M 0.001G
@@ VI Conclusion
In this paper, we have proposed SCBNN-OCM, a general framework that enhances variational inference-based Bayesian learning to target both ID and OOD calibration. To improve ID calibration, we have introduced a regularizer based on the calibration error, while OOD calibration is enhanced by means of data augmentation based on confidence minimization. In order to facilitate the synergistic use of calibration-aware regularization and OCM regularization, we have finally introduced a selective calibration strategy that rejects examples that are likely to have poor ID calibration performance. Numerical results have illustrated the challenges in ensuring both ID and OOD performance, as schemes designed for ID calibration may end up hurting OOD calibration, and vice versa. That said, thanks to selective calibration, SCBNN-OCM was observed to achieve the best ID and OOD performance as compared to the other benchmarks as long as one allows for a sufficiently large number of rejected samples. For instance, as compared to standard FNN, SCBNN-OCM achieves a $25\%$ improvement in accuracy, a $20\%$ drop in ECE, and a $50\%$ improvement in OOD detection probability, at the cost of reduced ID coverage rate of around $50\%$. Interesting directions for future work include extending the proposed framework to likelihood-free, simulation-based inference [56], to continual learning [57], and to large language models [13].
Supplementary Material for “Calibrating-Bayesian Learning via Regularization, Confidence Minimization, and Selective Inference”
VII ID Calibration Ablation
VII-A Impact of Calibration-Based Regularizer Weight
The impact of the calibration-based regularizer hyperparameter, $\lambda$ in (9) and (17), for ID calibration is analyzed in Fig. 11 under the same conditions as for Fig. 7 in the main text. The figure confirms that the best performance for CFNN and CBNN on CIFAR-100 (ID) and TinyImageNet (OOD) data sets is achieved for $\lambda=4$ and $\lambda=0.8$, respectively. CBNN scheme is seems to be more sensitive to the selection of hyperparameters $\lambda$, possibly because of the interplay with the choice of the prior distribution in the Bayesian learning objective (17). That said, CBNN outperforms BNN in terms of ECE for a wide range of values of hyperparameter $\lambda$.
Figure 11: ECE, accuracy and OOD detection probability as a function of hyperparameter $\lambda$ on CIFAR-100 (ID) and TinyImageNet (OOD) data set for FNN, BNN, CFNN and CBNN.
VIII OOD Detection Ablation
VIII-A Impact of OCM Regularizer Weight
Fig. 12 presents an ablation study for the OCM regularizer weight, $\gamma$ in (21) and (25), for the same setting as in the previous subsection. The ECE, test accuracy, and OOD detection probability are relatively stable when $\gamma>0.1$, which implies that our default choice of $\gamma=0.5$, also reported in [11], is well suited for all schemes.
Figure 12: ECE, accuracy and OOD detection probability as a function of hyperparameter $\gamma$ on CIFAR-100 (ID) and TinyImageNet (OOD) data set for FNN-OCM, BNN-OCM, CFNN-OCM and CBNN-OCM.
IX Selective Calibration Ablation
IX-A Impact of Selector Regularizer Weight
By varying the selector regularizer weight, $\eta$ in (33) and (39), Fig. 13 shows that the proposed Bayesian methods have a better performance in terms of all metrics as compared to SFNN, and are more robust to the choice of hyperparameter $\eta$. To maintain consistency with [12], we set $\eta=0.01$ as the default choice for $\eta$.
Figure 13: ECE, accuracy and OOD detection probability as a function of hyperparameter $\eta$ on CIFAR-100 (ID) and TinyImageNet (OOD) data set for SFNN, SBNN-OCM, and SCBNN-OCM with ID coverage rate as $40\%$.
X Backbone Study
In order to showcase that the gain of SCBNN-OCM is not restricted to a particular neural network architecture or to a specific ID/OOD data set, we present additional evaluations on a different backbone, namely ResNet-18, and on different data sets by considering CIFAR-10 as the ID and LSUN resized as the OOD data set. For all the experiments in this section, we use the same settings as the experimental results in the main text.
X-A Impact of the Calibration-Based Regularizer
The impact of the calibration-based regularization hyperparameter for ID calibration, $\lambda$ in (9) and (17), is analyzed in Fig. 14. The figure confirms that the best performance for CFNN and CBNN on CIFAR-10 (ID) and LSUN (OOD) data sets is achieved for $\lambda=10$ and $\lambda=4$, respectively. Also, increasing the value of $\lambda$ is seen to decrease the ECE for both CFNN and CBNN, with similar test accuracy.
X-B Impact of the OCM Regularizer
To investigate the impact of OCM regularizer on ID and OOD performance, Fig. 15 plots the ID and OOD performance by varying the OCM hyperparameter $\gamma$ in (21) and (25). Note that the deterioration of the ID performance, especially for CFNN-OCM, shows the side effects of introducing OCM, although it entails OOD detection probability improvement. Additionally, the proposed CBNN-OCM better balances the ID and OOD performance levels as compared to CFNN-OCM.
Figure 14: Accuracy versus ECE obtained by changing the hyperparameter $\lambda$ on CIFAR-10 data set for FNN, CFNN (benchmark), BNN, and CBNN (ours), with OOD detection probability indicated by the marker’s color.
Figure 15: Test accuracy versus OOD detection probability on CIFAR-10 (ID) and LSUN (OOD) for FNN-OCM (benchmark), CFNN-OCM, BNN-OCM, and CBNN-OCM (ours), with test ECE as the marker’s color. Note that hyperparameter $\gamma=0$ recovers the original schemes without OCM regularizer.
X-C Impact of Selective Calibration
To investigate the impact of selective calibration, Fig. 16 plots the ID and the OOD performance for different ID coverage rates ranging from $0.1$ to $1$. The figure shows that SCBNN-OCM outperforms SBNN-OCM in terms of ECE and OOD detection probability in different ID coverage rate regimes, while attaining similar test accuracy.
As compared to the Fig. 10 in the main text, SCBNN-OCM is seen to achieve better ID and OOD performance with relatively small reduction in terms of ID coverage rate. For instance, with an ID coverage rate of $0.7$, the ID accuracy is around $99\%$, the ECE is lower than $1$, and the OOD detection probability is nearly $1$.
Figure 16: Test accuracy, ECE, and OOD detection probability versus ID coverage rate on CIFAR-10 (ID) and LSUN (OOD) for SFNN (benchmark), SBNN-OCM and SCBNN-OCM (ours).
-D Non-Parametric Outlier Score Vectors
[…]
@@ X-C Impact of Selective Calibration
[…]
in which the coefficients $\alpha_{i}\geq 0$ for $i=1,...,|\mathcal{X}^{\text{tr}}|$, satisfying $\sum_{i=1}^{|\mathcal{X}^{\text{tr}}|}\alpha_{i}=1$, are optimized along with the offset parameter $\rho\in\mathbb{R}$ in order to separate the training data $\{z_{i}^{\text{tr}}\}_{i=1}^{|\mathcal{X}^{\text{tr}}|}$ from the origin in the feature space associated with the kernel $\kappa(\cdot)$ [59, Sec. 5].
@@
[…]
-E Experiment Details
@@
-E1 Architecture and Training Details for Calibration-Regularized Learning
@@
Hyperparameters: For fair comparison, we use the same training policy for both frequentist and Bayesian learning. Specifically, we use the SGD optimizer with momentum factor $0.9$ and train the model during $100$ epochs. We also decrease the learning rate by dividing its value by a factor of $5$ for every $30$ epochs with initial learning rate $0.1$ in a manner similar to [11]. The minibatch size used for single SGD update is set to $128$. For Bayesian learning, we set as prior $p(\theta)$ the Gaussian distribution that has zero-mean vector with diagonal covariance matrix having each element as $0.001$. The hyperparamter $\beta$ for the free energy (11) in the main text is set to $0.00035$. The ensemble size for Bayesian learning during training is set to $1$, while the ensemble size during testing is set to $20$. In a manner similar to [6], the calibration-based regularizer $\lambda$ in (9) and (17) in the main text is chosen as the value in set $\{0.2,0.4,0.6,0.8,1.0,2.0,3.0,...,10.0\}$ that achieves the lowest ECE, while preserving the accuracy drop no larger than $1.5\%$ as compared to a setting with $\lambda=0$. The corresponding accuracy and ECE are evaluated based on the validation data set $\mathcal{D}^{\text{val}}$, and the hyparparameter $\lambda$ for CFNN and for CBNN are chosen independently, which result in the values of $4$ and $0.8$ respectively in our experiments. With these choices, CFNN and CBNN can achieve lower ECE with acceptable accuracy drops. We use weighted MMCE regularizer for calibration-aware regularization term $\mathcal{E}(\theta|\mathcal{D}^{\text{tr}})$, which is defined as [4]
@@
[…]
@@
-E2 Architecture and Training Details for OOD Detection
@@
Hyperparameters: Following the original OCM paper [11], OOD confidence minimization (21) and (25) in the main text is carried out by fine-tuning based on the corresponding pre-trained models, e.g., CBNN-OCM is obtained via fine-tuning with the OCM-regularized training loss given the pre-trained CBNN. Uncertainty data set $\mathcal{D}^{\text{u}}$ is constructed by randomly choosing $6,000$ input data from TinyImageNet. During fine-tuning, SGD optimizer with momentum factor of $0.9$ is adopted during $10$ epochs, each consisting 282 iterations. The initial learning rate is set to $0.001$, and we update the learning rate for each iteration by following [62]. Specifically, during SGD training, for each epoch, we first sample 9,000 examples from $\mathcal{D}^{\text{tr}}$ and evaluate the ID-related training measures for each SGD update via sampling 32 examples (minibatch size being 32) without replacement among the 9,000 examples; for the OOD-related training measures, we sample 64 examples (minibatch size being 64) among $\mathcal{D}^{\text{u}}$ with replacement. The hyperparameter of OCM regularizer $\gamma$ in (21) and (25) in the main text is chosen as in reference [11], i.e, $\gamma=0.5$, which increases OOD detection probability nearly to $1$ for all OCM schemes. Other hyperparameters are the same as calibration-regularized learning described in the previous subsection.
@@
-E3 Architecture and Training Details for Selective Calibration
@@
[…]